OpenStudy (phebe):

HELPP IM DIEING LOLZZ 1. Find the pattern for the sequence and use it to show the next two terms. a. 51, 44, 37, 30… b. 5, 20, 80, 320…

OpenStudy (mathmath333):

1st one is AP with common difference -7=44-51=37-47..... 51-7=44,44-7=37,37-7=30,30-7=23 2nd is GP with common ratio 4,20/5=80/20=4....... 5,5*4=20,20*4=80,80*4=320,320*4=1280

OpenStudy (phebe):

ok thanx soo much

OpenStudy (mathmath333):

welcome